### v1

@@ 6 Application to TB study
In this section, we assess the efficiency of the EGDL-Parallel and EGDL-Series frameworks by comparing their TB incidence forecasting performance across Japan’s 47 prefectures with state-of-the-art architectures. To evaluate the robustness and generalizability of the EGDL-based approaches, we employ rolling window forecast horizons of 12 months, 9 months, 6 months, and 3 months, performing multi-step ahead forecasts for each respective horizon. For the 12-month evaluation period, the models are trained using incidence data recorded from January 1998 to December 2014, while the remaining observations are reserved for testing. Similarly, for the 9-month, 6-month, and 3-month horizons, training data spans from January 1998 to March 2015, June 2015, and September 2015, respectively, with the remaining data used for evaluation. Since EGDL is an integrated approach of the networked SIR model with a saturated incidence rate and modern deep learning-based time series forecasting methods, therefore, we first delve into the parameter estimation procedure for the former parts.
6.1 TB data characteristics
@@
The monthly total number of active TB cases recorded across the 47 prefectures of Japan is devoid of any missing observations and demonstrates rapid variation in incidence cases. The average monthly incidence ranges from 9.49 to 294.84 cases across different prefectures. We checked several global characteristics for spatiotemporal TB datasets and reported the numerical results in Table A4 of Appendix C. The results of the statistical tests reveal that the TB incidence time series for all 47 prefectures of Japan are non-stationary and exhibit long-term dependency, as indicated by Hurst exponents greater than 0.50. Additionally, most series display non-linear patterns with evident quarterly and annual seasonality. Furthermore, we explore the spatial relationships among prefectures by analyzing pairwise correlations in TB incidence. Fig. A2(b) (Appendix C) presents a heatmap of these correlations, showing stronger associations between geographically proximate prefectures. For example, prefecture 10 (Gunma) demonstrates high positive correlations with its neighboring regions, while prefecture 47 (Okinawa) exhibits weaker correlations due to its geographic isolation. Diagonal entries in the heatmap represent self-correlations, which equals 1. Some distant prefectures (e.g., prefectures 3 and 20) exhibit moderate to strong correlations, likely due to shared non-local influences. These findings underscore the simultaneous spatial and temporal dependencies inherent in TB incidence cases in Japan, emphasizing the importance of considering both dimensions when modeling and forecasting disease dynamics.
6.2 Parameter Calibration
@@
To numerically solve the networked SIR model with saturated incidence rate and graph Laplacian diffusion Eq. (11), we transform it into a system of ODEs using Eq. (4). This process requires the Laplacian matrix used in our study to solve the model numerically. The Laplacian matrix $L$ is defined by $L=D-A,$ where $D$ is the degree matrix and $A$ is the adjacency matrix. The degree matrix $D$ is a diagonal matrix with the entries $D_{ii}$, the degree of node $i$, representing the number of edges incident to node $i$ i.e., $D_{ii}=\text{deg}(v_{i})$. The adjacency matrix $A$ with entries $A_{ij}$ indicates the link between nodes $i$ and $j$ as follows:
@@
$A_{ij}=\begin{cases}1,&\text{if there is a link between nodes }i\text{ and }j\\ 0,&\text{otherwise}\end{cases}$ (13)
@@
Thus, the diagonal elements of $L$ are the degrees of the nodes, while the off-diagonal elements $L_{ij}$ are -1 if nodes $i$ and $j$ are connected by an edge, and 0 otherwise:
$L_{ij}=\begin{cases}\text{deg}(v_{i}),&\text{if }i=j\\ -1,&\text{if }i\neq j\text{ and }(i,j)\in\mathcal{E}\\ 0,&\text{if }i\neq j\text{ and }(i,j)\notin\mathcal{E}.\end{cases}$ (14)
@@
To construct the Laplacian matrix across different prefectures of Japan, we assume a connection between two prefectures if they share a common boundary. For the prefectures of Hokkaido, Tokushima, Kagawa, Ehime, and Okinawa, which are geographically separated from the mainland (see Fig. 1(A)), we establish connections to the nearest mainland prefecture. The structure of these connections is illustrated in Fig. A2(a) of Appendix C. Now, for each node $k\in V$, we rewrite the model in Eq. (11) as follows:
@@
Using the Laplacian matrix (see Fig. A2(a) of Appendix C), we numerically solve the system Eq. (15), assuming that prefecture 20 (Nagano) serves as the source of infection. Additionally, to support theoretical findings, we consider an initial total population of 47,000 distributed evenly across the 47 prefectures, with each prefecture having a population of 1,000. The initial conditions are defined as follows:
@@
Now, we calculate the basic reproduction number, $\mathcal{R}_{0}=0.385$, that is less than unity, according to Theorem 1 and the disease-free equilibrium $(\Lambda/\mu,0)=(100.0,0)$ point is globally asymptotically stable (see Fig A3(a-b) of Appendix C.1). In Fig A3(a), all $S$ solutions converge to $\Lambda/\mu$ regardless of the node, similarly, in Fig A3(b), all $I$ solutions converge to 00 regardless of the node. These results can be similarly demonstrated, independent of the initial conditions. For the endemic equilibrium, we use all the parameters from Eq. (17) except $\beta$ and $\sigma$ which are set to $\beta=0.001$ and $\sigma=10^{-5}$. Direct calculation gives: $\mathcal{R}_{0}=3.85,S^{*}=876.67,I^{*}=7.74$. Since $\mathcal{R}_{0}=3.85$ is greater than unity, then by Theorem 2 endemic equilibrium $(S^{*},I^{*})=(876.67,7.74)$ point is globally asymptotically stable as shown in Fig A3(c-d). In Fig A3(c), all $S$ solutions converge to $S^{*}$ regardless of the node; similarly, in Fig A3(d), all $I$ solutions converge to $I^{*}$ regardless of the node. The saturation parameter $\alpha$ significantly influences the solution of the system, as illustrated in Fig. A4 of Appendix C.1. For this analysis, we use all the parameters associated with the endemic equilibrium case, except for $\alpha$, which is varied between 0.1 and 1.0. At lower values of $\alpha$, the contact rate increases, resulting in a higher peak of infections. As $\alpha$ approaches 1.0, the contact rate saturates, and the infection profile becomes more uniform over time.
6.3 Sensitivity Analsyis
@@
In this part, we estimate the posterior distributions of the networked SIR model parameters using the Markov-Chain Monte-Carlo (MCMC) approach on the Laplacian network (depicted in Fig. A2(a) of Appendix C) via the PyMC Python library [53]. To get posterior distributions of the unknown parameters, the state variables $\hat{S}_{i}(t),\hat{I}_{i}(t)$ are observed at $T$ certain times $t_{1},t_{2},\dots,t_{T}$. Let, $\theta=\{\alpha,\beta,\gamma,\sigma,\mu\}$ be the unknown parameter and $D_{i}=\{\hat{S}_{i},\hat{I}_{i}\}$ be observed data at node $i$. It is not always possible to find sample $\theta$ from the posterior probability distribution $\textbf{P}(\theta\,|\,D_{i})$. This is where the iterative MCMC algorithm plays a crucial role. It generates a new vector parameter at step $\hat{n}$ ($\theta^{\hat{n}}$), from the posterior distribution, using the previous vector parameter ($\theta^{\hat{n}-1}$), with initial guess $\theta^{0}$. The chain or sample-path run $\hat{n}$ steps until it reaches its stationary distribution, where $\hat{n}$ is a sufficiently large number, and fills the space of the target un-normalized posterior distribution. For implementing Bayesian inference with MCMC, we used the No-U-Turn Sampler (NUTS) [54]. The NUTS approach automatically selects these parameters, making it a tuning-free sampling algorithm that performs similar or better than other MCMC algorithms [54]. Before performing inference, we estimate the parameter set $\theta$, denoted as $\theta_{ls}\equiv\{\alpha_{ls},\beta_{ls},\gamma_{ls},\sigma_{ls},\mu_{ls}\}$, using the ordinary least squares (OLS) method with TB dataset from 47 prefectures in Japan. Based on these estimates, we define the prior distribution information for the parameters as follows:
@@
where $\eta$ is the likelihood distribution. We selected the ‘TruncatedNormal’ distribution to ensure the non-negativity of the parameters, aligning with epidemiological considerations. For each MCMC step $\hat{n}$, NUTS choose the new parameter $\theta^{\hat{n}}$ in such a way that the set of parameters minimizes the previous deviation. Each time $t$, integration with its model parameters produces one Monte Carlo sample, and iteratively, MCMC will create the full set of required samples. The summary of the MCMC results is provided in Table 1. Now, using the values of the estimated parameters for the networked SIR model, we generate the predictions for the infected curve which will be further used in the EGDL framework.
@@
6.4 Forecasting Performance Evaluation
@@
This section first examines the causal relationship between TB incidence cases and the infected curve of the networked SIR model (see Appendix C.2 for details). The results show a strong causal relationship between TB incidence cases and the infection dynamic curve. This finding supports the integration of historical incidence data and the networked SIR model’s infection curve in the EGDL architectures, leveraging the complementary strengths of both approaches to improve TB incidence forecasting. Furthermore, to evaluate the performance of the forecasting approaches, we utilize four key performance indicators, namely Symmetric Mean Absolute Percent Error (SMAPE), Mean Absolute Error (MAE), Mean Absolute Scaled Error (MASE), and Root Mean Squared Error (RMSE) [55] (details in Appendix C.3). By convention, the lower values of these metrics indicate better performance. In this study, we report the average performance metrics computed across the 47 prefectures of Japan for various forecast horizons. Using these four performance measures, we evaluate the performance of the EGDL-Parallel and EGDL-Series frameworks for forecasting TB incidence cases. While these frameworks can be flexibly integrated with any forecasting model, our analysis employs state-of-the-art global architectures, including Transformers [39], NBeats [41], NHits [42], and TCN [40]. These models are specifically designed for time series forecasting and are well-suited to multivariate setups, which is challenging for many traditional statistical methods. By integrating the EGDL-Parallel and EGDL-Series frameworks with the Transformers architecture, we develop the epidemic-guided parallel Transformers (EGP-Transformers) and epidemic-guided series Transformers (EGS-Transformers), respectively. Similarly, integrating NBeats with these frameworks results in the epidemic-guided parallel NBeats (EGP-NBeats) and epidemic-guided series NBeats (EGS-NBeats), respectively. Following the same methodology, combining EGDL-Parallel and EGDL-Series with NHits and TCN generates EGP-NHits, EGS-NHits, EGP-TCN, and EGS-TCN models, respectively. These newly developed forecasting models effectively learn temporal patterns from lagged values while incorporating epidemic dynamics through the estimated infection curves by taking the mean of the model parameters from Table 1 using Bayesian inference with the MCMC approach from the networked SIR model. This integration ensures that epidemic principles, including spatial interactions, guide the forecasters to predict realistic future disease trends. To assess the impact of incorporating these epidemic principles, we compare the performance of the epidemic-guided models with their corresponding baseline architectures. Additionally, we include three spatiotemporal forecasting techniques, namely GSTAR [46], GpGp [47], and STGCN [48], which can inherently model spatial and temporal dependencies in the disease incidence cases.
Tables 2 and 3 summarize the forecasting performance of the EGDL-Parallel and EGDL-Series frameworks alongside baseline models for predicting TB incidence in Japan’s 47 prefectures across various forecast horizons. As shown in Table 2, the EGP-NHits model consistently delivers superior performance in most forecasting tasks. The integration of epidemic knowledge, specifically the networked SIR model’s estimated infection curve, enables the EGP-NHits framework to produce reliable forecasts, significantly enhancing its accuracy compared to its baseline architecture. Similarly, the EGP-Transformers model demonstrates notably higher accuracy than the individual Transformers and networked SIR models. The EGP-NBeats and EGP-TCN architectures also show marked improvements over their respective baselines, further justifying the advantage of the combined forecasting techniques. Additionally, Table 3 reveals the consistent superiority of the EGDL-Series frameworks over competing models across various forecast horizons. For the 12-month forecast period, the EGS-NBeats, EGS-TCN, and EGS-NHits frameworks demonstrate the best performance across multiple indicators. For the 9-month horizon, the EGS-TCN model provides the most accurate out-of-sample predictions, followed closely by the EGS-NHits architecture. Similar trends are evident for the 6-month forecast horizon, where EGS-NHits and EGS-TCN outperform all competing models. In short-term forecasting over 3 months, the spatiotemporal GSTAR and GpGp models offer competitive performance in terms of MAE and RMSE, while the EGS-Transformers and NHits models deliver the lowest MASE and SMAPE values, respectively. Overall, the combination of deterministic and stochastic components, achieved through residual remodeling in the EGDL-Series frameworks, significantly improves the performance of individual deep learning models and the networked SIR model across most forecast horizons. Furthermore, we employ the non-parametric multiple comparisons with the best (MCB) test to validate the statistical significance of the performance improvements [56]. The results of the MCB test (see Appendix C.4) showcase that all forms of EGDL architectures improve the performance of the corresponding individual models due to their unique merit in integrating epidemiological principles, spatial dynamics, and temporal interactions of TB datasets. More particularly, the performance of the data-centric approaches improved by employing the residual remodeling mechanism of EGDL-Series frameworks. Overall, the EGDL-Series architectures are better suited for medium-term (9-month and 12-month) forecasting tasks, whereas the EGDL-Parallel framework provides more accurate short-term (up to semi-annual) forecasts of TB incidence. Additionally, we quantify the uncertainties associated with the EGDL frameworks (Fig. 2) using the conformal prediction approach (see Appendix C.5). The overall empirical results and prediction intervals provide significant evidence for the practical applicability of EGDL architectures in designing timely and targeted public health intervention strategies.
@@
Table 2: Forecast performance of spatiotemporal models, temporal deep learners, and Epidemic-Guided Deep Learning (EGDL)-Parallel approach for different forecast horizons. The best results are highlighted.
@@
GSTAR GpGp STGCN Networked Trans- NBeats NHits TCN EGP-Trans- EGP-NBeats EGP-NHits EGP-TCN
SIR formers formers
@@
Table 3: Forecast performance of spatiotemporal models, temporal deep learners, and Epidemic-Guided Deep Learning (EGDL)-Series approach for different forecast horizons. The best results are highlighted.
@@
GSTAR GpGp STGCN Networked Trans- NBeats NHits TCN EGS-Trans- EGS-NBeats EGS-NHits EGS-TCN
SIR formers formers

### v2

@@ 6 Application to TB study
In this section, we assess the efficiency of the EGDL-Parallel and EGDL-Series frameworks across two geographically and epidemiologically distinct settings, Japan and China. For both countries, we compare our models against state-of-the-art forecasting architectures using rolling windows over 12, 9, 6, and 3-month horizons for evaluating the accuracy and generalizability of the EGDL frameworks across diverse temporal contexts and population structures. For Japan, the experiments are conducted across all 47 prefectures. The models are trained using monthly TB incidence data from January 1998 to December 2014 for the 12-month horizon, while the subsequent months are used for testing. Similarly, for the 9-month, 6-month, and 3-month horizons, the training periods extend until March 2015, June 2015, and September 2015, respectively, with the remaining data reserved for evaluation. In the case of China, we retain a consistent experimental setup to ensure a fair comparison. The models are trained on monthly TB incidence data of 31 provinces from January 2014 to December 2017 for the 12-month horizon, with the remaining months of 2018 used for testing. For the 9-month and 6-month forecasts, training extends up to March 2018 and June 2018, respectively. For the 3-month horizon, data up to September 2018 is used for training, while the final quarter of 2018 is held out for testing. Since EGDL integrates MN-SIR model with deep learning-based forecasting techniques, we begin by detailing the data characteristics and parameter estimation process of the mechanistic component before analyzing the overall forecasting performance across both countries.
6.1 TB Data Characteristics of Japan and China
(A) Laplacian Matrix (B) Correlation Plot
Refer to caption Refer to caption
Figure 4: (A) The Laplacian matrix corresponding to the network depicted in Fig. 1(A), representing the connectivity and structure of the nodes in the system. This matrix captures the relationships between neighboring nodes and is essential for analyzing diffusion processes and other network dynamics. (B) Pairwise correlation of TB incidence cases recorded at 47 prefectures of Japan.
@@ 6.1 TB Data Characteristics of Japan and China
The TB incidence datasets for Japan and China, though collected from distinct geographical and epidemiological settings, both exhibit strong spatiotemporal dependencies that are crucial for accurate forecasting. In Japan, monthly TB case counts across 47 prefectures are devoid of missing observations and exhibit rapid variations, with average monthly incidences ranging from 9.49 to 294.84. We analyzed several global characteristics for spatiotemporal TB datasets. The results of the statistical analysis, reported in Table A8 of Appendix C, reveal that the TB incidence time series for all 47 prefectures of Japan are non-stationary and exhibit long-term dependency, as indicated by Hurst exponents greater than 0.50. Additionally, most series display non-linear patterns with quarterly and annual seasonality. Furthermore, we explore the spatial relationships among prefectures by analyzing pairwise correlations in TB incidence. Fig. 4(B) presents a heatmap of these correlations, showing stronger associations between geographically proximate prefectures, for instance prefecture 10 (Gunma) and its neighboring regions, while isolated prefectures like 47 (Okinawa) exhibit weaker correlations. Diagonal entries in the heatmap represent self-correlations, which are equal to 1. A similar spatiotemporal structure is also observed in mainland China’s TB incidence cases. This dataset, free from missing observations, also exhibits considerable regional variations, with the average monthly cases ranging from 243.67 to 4,877.45. Statistical analysis of mainland China’s TB incidence datasets, summarized in Table A9 of Appendix C, reveals that all series exhibit long-term memory, are predominantly non-stationary, follow approximately linear trends, and demonstrate annual seasonality. Correlation heatmaps, Fig. 5(B), indicate that geographically adjacent provinces, such as 3 (Hebei), 4 (Shanxi), and 5 (Inner Mongolia), tend to have stronger correlations, while remote provinces such as 1 (Beijing) and 26 (Tibet) show weaker associations. Notably, both TB incidence data from Japan and China reveal moderate to strong correlations even between non-adjacent regions, likely due to shared non-local drivers. These findings collectively validate the presence of both spatial and temporal dependencies in TB incidence across Japan and China, emphasizing the importance of jointly modeling these dynamics to improve forecasting accuracy.
Figure 5: (A) The Laplacian matrix corresponding to the network depicted in Fig. 2(A), representing the connectivity and structure of the nodes in the system. This matrix captures the relationships between neighboring provinces in mainland China and is essential for analyzing diffusion processes and other network dynamics. (B) Pairwise correlation of TB incidence cases recorded at 31 provinces of mainland China.
6.2 Results on Japan’s Prefecture-specific Active TB Cases
In this section, we conduct an in-depth analysis of the EGDL frameworks for forecasting monthly active TB cases across 47 prefectures of Japan. We begin by estimating the key epidemiological parameters that govern infection dynamics within the MN-SIR model, followed by a comprehensive evaluation of the forecasting performance of the proposed frameworks.
6.2.1 Parameter Calibration
@@
To numerically solve the MN-SIR model, as in Eq. (12), we transform it into a system of ODEs (15) using the definition (5). This process requires the Laplacian matrix used in our study to solve the model numerically. The Laplacian matrix $L$ is defined by $L=D-A,$ where $D$ is the degree matrix and $A$ is the adjacency matrix. The degree matrix $D$ is a diagonal matrix with the entries $D_{ii}$, the degree of node $i$, representing the number of edges incident to node $i$, i.e., $D_{ii}=\text{deg}(v_{i})$. The adjacency matrix $A$ with entries $A_{ij}$ indicates the link between nodes $i$ and $j$ as follows:
@@
$A_{ij}=\begin{cases}1,&\text{if there is a link between nodes }i\text{ and }j\\ 0,&\text{otherwise.}\end{cases}$ (13)
@@
Thus, the diagonal elements of $L$ are the degrees of the nodes, while the off-diagonal elements $L_{ij}$ are 1 if nodes $i$ and $j$ are connected by an edge, and 0 otherwise:
$L_{ij}=\begin{cases}-\text{deg}(v_{i}),&\text{if }i=j\\ 1,&\text{if }i\neq j\text{ and }(i,j)\in\mathcal{E}\\ 0,&\text{if }i\neq j\text{ and }(i,j)\notin\mathcal{E}.\end{cases}$ (14)
@@
To construct the Laplacian matrix across different prefectures of Japan, we assume a connection between two prefectures if they share a common boundary. For the prefectures of Hokkaido, Tokushima, Kagawa, Ehime, and Okinawa, which are geographically separated from the mainland (see Fig. 1(A)), we establish connections to the nearest mainland prefecture. The structure of these connections is illustrated in Fig. 4(A). Now, for each node $k\in V$, we rewrite the model in Eq. (12) as follows:
@@
Using the Laplacian matrix (see Fig. 4(A)), we numerically solve the system Eq. (15), assuming that prefecture 20 (Nagano) serves as the source of infection. Additionally, to support theoretical findings, we consider an initial total population of 47,000 distributed evenly across the 47 prefectures, with each prefecture having a population of 1,000. The initial conditions are defined as follows:
@@
Now, we calculate the basic reproduction number, $\mathcal{R}_{0}=0.385$, that is less than unity, according to Theorem 1 and the disease-free equilibrium $(\Lambda/\mu,0)=(1000.0,0)$ point is globally asymptotically stable (see Fig A1(a-b) of Appendix C.1). In Fig A1(a), all $S$ solutions converge to $\Lambda/\mu$ regardless of the node, similarly, in Fig A1(b), all $I$ solutions converge to 00 regardless of the node. These results can be similarly demonstrated, independent of the initial conditions. For the endemic equilibrium, we use all the parameters from Eq. (17) except $\beta$ and $\sigma$ which are set to $\beta=0.001$ and $\sigma=10^{-5}$. Direct calculation gives: $\mathcal{R}_{0}=3.85,S^{*}=876.67,\text{ and }I^{*}=4.74$. Since $\mathcal{R}_{0}=3.85$ is greater than unity, then by Theorem 2 endemic equilibrium $(S^{*},I^{*})=(876.67,7.74)$ point is globally asymptotically stable as shown in Fig A1(c-d). In Fig A1(c), all $S$ solutions converge to $S^{*}$ regardless of the node; similarly, in Fig A1(d), all $I$ solutions converge to $I^{*}$ regardless of the node. The saturation parameter $\alpha$ significantly influences the solution of the system, as illustrated in Fig. A2 of Appendix C.1. For this analysis, we use all the parameters associated with the endemic equilibrium case, except for $\alpha$, which is varied between 0.1 and 1.0. At lower values of $\alpha$, the contact rate increases, resulting in a higher peak of infections. As $\alpha$ approaches 1.0, the contact rate saturates, and the infection profile becomes more uniform over time.
6.2.2 Sensitivity Analsyis
@@
In this part, we estimate the posterior distributions of the MN-SIR model parameters using the Markov-Chain Monte-Carlo (MCMC) approach on the Laplacian network (depicted in Fig. 4(A)) via the PyMC Python library [58]. To get posterior distributions of the unknown parameters, the state variables $\hat{S}_{i}(t)$ and $\hat{I}_{i}(t)$ are observed at $T$ certain times $(t_{1},t_{2},\dots,t_{T})$. Let, $\theta=\{\alpha,\beta,\gamma,\sigma,\mu\}$ be the unknown parameter and $D_{i}=\{\hat{S}_{i},\hat{I}_{i}\}$ be observed data at node $i$. It is not always possible to find sample $\theta$ from the posterior probability distribution $\textbf{P}(\theta\,|\,D_{i})$. This is where the iterative MCMC algorithm plays a crucial role. It generates a new vector parameter at step $\hat{n}$ ($\theta^{\hat{n}}$), from the posterior distribution, using the previous vector parameter ($\theta^{\hat{n}-1}$), with initial guess $\theta^{0}$. The chain or sample-path run $\hat{n}$ steps until it reaches its stationary distribution, where $\hat{n}$ is a sufficiently large number, and fills the space of the target un-normalized posterior distribution. For implementing Bayesian inference with MCMC, we utilized the No-U-Turn Sampler (NUTS) that adaptively tunes its parameters and performs competitively with other sampling methods. [59]. In this study, the NUTS sampler was configured with two chains, each comprising 3,000 tuning (warm-up) and 3,000 sampling iterations, ensuring convergence to a stationary distribution and accurate parameter estimates. On a single core of an Intel i7 9th generation processor, each chain completed in approximately 2.5 minutes, demonstrating the feasibility of our approach for moderate-scale applications on standard hardware. For larger datasets or real-time deployment, this process could be accelerated using parallel computing, hardware acceleration, or approximate inference techniques.
To initialize the Bayesian inference process, we first estimate the parameter set $\theta$, denoted as $\theta_{ls}\equiv\{\alpha_{ls},\beta_{ls},\gamma_{ls},\sigma_{ls},\mu_{ls}\}$, using the ordinary least squares (OLS) method with TB dataset from 47 prefectures in Japan. Based on these estimates, we define the prior distribution information for the parameters as follows:
@@
where $\eta$ is the likelihood distribution. We selected the ‘TruncatedNormal’ distribution to ensure the non-negativity of the parameters, aligning with epidemiological considerations. For each MCMC step $\hat{n}$, NUTS chooses the new parameter $\theta^{\hat{n}}$ in such a way that the set of parameters minimizes the previous deviation. At each time $t$, integration with its model parameters produces one Monte Carlo sample, and iteratively, MCMC will create the full set of required samples. Table 1 summarizes the results of the MCMC sampling, conducted using the NUTS approach with 3000 samples across 2 chains, based on the prior specifications given in Eq. (18). The reported statistics include the posterior mean (Mean), standard deviation (SD), highest density intervals (HDI) at 3% and 97%, Monte Carlo standard errors (MCSE) for both the mean and SD, effective sample sizes (ESS) for both bulk and tail estimates, and the potential scale reduction factor (R_hat), which assesses convergence across chains. The parameter, $\alpha$, has a posterior mean of 0.04112 with a narrow standard deviation SD=0.00116, indicating good confidence in its estimate. The 94% HDI range, given by HDI = [0.03888, 0.04324], also confirms this precision. The convergence diagnostics R_hat =1.00793 and bulk ESS = 64.44 suggest moderate sampling efficiency and acceptable convergence. The parameters $\beta,\gamma$ can be interpreted similarly. The parameter $\sigma$ and the parameter $\mu$ are both small in magnitude. They are estimated with high precision and sampling efficiency, as indicated by large ESS values (bulk ESS) and R_hat $\approx 1$, confirming excellent convergence. Lastly, $\eta$ also exhibits very good convergence. Now, using the values of the estimated parameters for the MN-SIR model, we generate the predictions for the infected curve, which will be further used in the EGDL framework.
@@
6.2.3 Forecasting Performance Evaluation
@@
To evaluate the performance of the forecasting approaches, we utilize four key performance indicators, namely Symmetric Mean Absolute Percent Error (SMAPE), Mean Absolute Error (MAE), Mean Absolute Scaled Error (MASE), and Root Mean Squared Error (RMSE) [36] (details in Appendix C.2). By definition, the lower values of these metrics indicate better performance. In this evaluation, we report the average performance metrics and corresponding standard deviations (SDs) computed across 47 prefectures of Japan for various forecast horizons. To ensure robustness and consistency, the forecasting models are trained over 5 independent runs, and the reported results reflect the aggregated performance across these repetitions. While the EGDL-Parallel and EGDL-Series frameworks can be flexibly integrated with any forecasting model, our analysis employs state-of-the-art global architectures, including Transformers, NBeats, NHits, and TCN. These models are specifically designed for time series forecasting and are well-suited to multivariate setups, which is a key challenge for many traditional statistical methods [60]. By integrating the EGDL-Parallel and EGDL-Series frameworks with the Transformers architecture, we develop the epidemic-guided parallel Transformers (EGP-Transformers) and epidemic-guided series Transformers (EGS-Transformers), respectively. Similarly, integrating NBeats with these frameworks results in the epidemic-guided parallel NBeats (EGP-NBeats) and epidemic-guided series NBeats (EGS-NBeats), respectively. Following the same methodology, combining EGDL-Parallel and EGDL-Series with NHits and TCN generates EGP-NHits, EGS-NHits, EGP-TCN, and EGS-TCN models, respectively. These newly developed forecasting models effectively learn temporal patterns from lagged values while incorporating epidemic dynamics via the infection curves estimated from the MN-SIR model, using the mean of the parameters reported in Table 1. This integration ensures that epidemic principles, including spatial interactions, guide the forecasters to predict realistic future disease trends. To assess the impact of incorporating these epidemic principles, we compare the performance of the epidemic-guided models with their corresponding baseline architectures. Additionally, we include three spatiotemporal forecasting techniques, namely GSTAR [51], GpGp [52], and STGCN [53], which can inherently model spatial and temporal dependencies in the disease incidence cases.
@@
Table 2: Mean forecast performance along with standard deviations (SDs) of spatiotemporal models, temporal deep learners, and Epidemic-Guided Deep Learning (EGDL)-Parallel approach for different forecast horizons on TB incidence data of Japan. The best results are highlighted.
@@
GSTAR GpGp STGCN MN-SIR Trans- NBeats NHits TCN EGP-Tran- EGP- EGP- EGP-
formers sformers NBeats NHits TCN
@@
(3.88) (3.47) (16.51) (6.75) (27.88) (4.40) (4.51) (9.48) (28.04) (4.30) (4.03) (4.56)
Table 3: Mean forecast performance along with standard deviations (SDs) of spatiotemporal models, temporal deep learners, and Epidemic-Guided Deep Learning (EGDL)-Series approach for different forecast horizons on TB incidence data of Japan. The best results are highlighted.
@@
GSTAR GpGp STGCN MN-SIR Trans- NBeats NHits TCN EGS-Tran- EGS- EGS- EGS-
formers sformers NBeats NHits TCN